OpenStudy (anonymous):

Find the inverse of the function. f(x) = 7x + 2

OpenStudy (anonymous):

just solve for x

OpenStudy (anonymous):

=7x+2?

OpenStudy (anonymous):

well if you are asked to change a function into inverse. Just replace the x with y or viceversa. then, solve for the y. y = 7x + 2 x = 7y + 2 x-2 = 7y y = (x-2 )/ 7 <-- answer

OpenStudy (anonymous):

f(x)=y=7x+2 1 x = (1/7)*y-2/7 sfx 2 y = (1/7)*x-2/7 swap x for y

OpenStudy (anonymous):

noo i want the inverse of the function

OpenStudy (anonymous):

so x+2/7

OpenStudy (anonymous):

y = (1/7)*x-2/7 is inverse

OpenStudy (anonymous):

^that is basically, y = (x-2)/7

OpenStudy (anonymous):

ok thank you i have a few more if u guys don't mind helping :/ &&& thank you

OpenStudy (anonymous):

What is the aversion to leaving the inverse as a function of y? It is the same thing and is easier to differentiate.